OpenStudy (anonymous):

Which set does -2 belong to?

OpenStudy (anonymous):

infinitely many sets

OpenStudy (anonymous):

do you have more information?

OpenStudy (anonymous):

I have no further info, sadly. It just asks "Name the set or sets to which each number belongs" And the number is -2

OpenStudy (anonymous):

it is an integer, a whole number, a real number, a natural number and a rational number

jimthompson5910 (jim_thompson5910):

the set of whole numbers is {0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, ...} the set of natural numbers is {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, ...}

OpenStudy (anonymous):

Thank you! The last one asks for 320/2 in a square root symbol. Sorry, my math book didn't go over this and I'm out of school for awhile and can't figure this out.

jimthompson5910 (jim_thompson5910):

320/2 = ???

OpenStudy (anonymous):

12.64911064 I believe.

jimthompson5910 (jim_thompson5910):

ok you're saying \[\large \sqrt{\frac{320}{2}} = \sqrt{160} \approx 12.64911064\] and you are correct

OpenStudy (anonymous):

So then what set(s) would that be a part of?

jimthompson5910 (jim_thompson5910):

since you get some decimal number that goes on forever without repeating, this means that \(\large \sqrt{\frac{320}{2}}\) is irrational

jimthompson5910 (jim_thompson5910):

it means that you cannot represent that as a fraction of two whole numbers

OpenStudy (anonymous):

Okay, this makes a lot of sense and thank you. My notes I was sent did not explain any of this. I don't understand my school. Oh well.

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

it belongs to selt of real numbers